The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Bryan C. Dunne Co-Editors Eva K. Grebel and You-Hua Chu

                                 mcnews@astro.uiuc.edu
                  http://www.astro.uiuc.edu/projects/mcnews/MCNews.php

No. 71 July 30th, 2003

Contents

9 Abstracts of Refereed Papers  2
1 Abstract of a Non-Refereed Paper 8
Recent astro-ph Listings  8

Letter From The Editor

Dear Colleagues,

The Magellanic Clouds Working Group is pleased to bring you issue 71 of the Magellanic Clouds Newsletter. This month, we present 10 submitted abstracts and a collection of recent astro-ph listings of interest to Magellanic Clouds researchers. In addition, we’d like to call your attention to this month’s Hubble Heritage Project release (http://heritage.stsci.edu), which features a fabulous image of the supernova remnant N49 in the Large Magellanic Cloud.

Best Wishes, Bryan Dunne Editor, MC News

Abstracts of Refereed Papers

WR bubbles and HeII emission

Y. Nazé (1), G. Rauw (1), J. Manfroid (1), Y.-H. Chu (2), and J.-M. Vreux (1)

(1) Institut d’Astrophysique et de Géophysique, Université de Liège, Liège, Belgium

(2) Astronomy Department, University of Illinois at Urbana-Champaign, Urbana, IL, USA

We present the very first high quality images of the HeII λ4686 emission in three high excitation nebulae of the Magellanic Clouds. A fourth high excitation nebula, situated around the WR star BAT99-2, was analysed in a previous letter. Using VLT FORS data, we investigate the morphology of the ring nebulae around the early-type WN stars BAT99-49 & AB7. We derive the total HeII fluxes for each object and compare them with the most recent theoretical WR models. Whilst the ionization of the nebula around BAT99-49 can be explained by a WN star of temperature 90-100 kK, we find that the HeII emission measure of the nebula associated with AB7 requires an He^{+} ionizing flux larger than predicted for the hottest WN model available. Using Hα, [OIII] and HeI λ5876 images along with long-slit spectroscopy, we investigate the physical properties of these ring nebulae and find only moderate chemical enrichment.

We also surveyed seven other LMC WR stars but we failed to detect any HeII emission. This holds also true for BAT99-9 which had been proposed to excite an HeII nebula. Four of these surveyed stars are surrounded by a ring nebula, and we use the FORS data to derive their chemical composition: the nebula around BAT99-11 shows a N/O ratio and an oxygen abundance slightly lower than the LMC values, while the nebula around BAT99-134 presents moderate chemical enrichment similar to the one seen near BAT99-2, 49 and AB7. Comparing the WR stars of the LMC, BAT99-2 and 49 appear rather unique since similar stars do not reveal high excitation features.

The third high excitation nebula presented in this paper, N44C, does not harbor stars hotter than mid-O main sequence stars. It was suggested to be a fossil X-ray nebula ionized by the transient LMC X-5. Our observations of N44C reveal no substantial changes in the excitation compared to previous results reported in the literature. Therefore, we conclude that either the recombination timescale of the X-ray nebula has been underestimated or that the excitation of the nebula is produced by another, yet unknown, mechanism. Comments: accepted by A&A, also available as astro-ph/0306084 WWW: http://vela.astro.ulg.ac.be/Preprints/P81/index.html e-mail: naze@astro.ulg.ac.be

On the Structure and Kinematics of Nebulae around LBVs and LBV Candidates in the LMC

Kerstin Weis

ITA, University of Heidelberg, Heidelberg, Germany; University of Minnesota, Astronomy Department, MN, USA; MPIfR, Bonn, Germany

We present a detailed analysis of the morphology and kinematics of nebulae around LBVs and LBV candidates in the Large Magellanic Cloud. HST images and high-resolution Echelle Spectra were used to determine the size, shape, brightness, and expansion velocities of the LBV nebulae around R 127, R 143, and S 61. For S Dor, R 71, R 99, and R 84 we discuss the possible presence of nebular emission, and derive upper limits for the size and lower limits on the expansion velocities of possible nebulae. Including earlier results for the LBV candidates S 119 and Sk-69 279 we find that in general the nebulae around LBVs in the LMC are comparable in size to those found in the Milky Way. The expansion velocities of the LMC nebulae, however, are significantly lower—by about a factor of 3 to 4—than those of galactic nebulae of comparable size. Galactic and LMC nebulae show about the same diversity of morphologies, but only in the LMC do we find nebulae with outflow. Bipolarity—at least to some degree—is found in nebulae in the LMC as well as in the Milky Way, and manifests a much more general feature among LBV nebulae than previously known

Comments: accepted in A&A WWW: http://www.astro.rub.de/kweis/publications.html e-mail: kweis@astro.rub.de

The physical properties of the compact star forming region N88A

and its neighbourhood in the SMC

I. Excitation and abundances

G. Testor (1), J.L. Lemaire (1 & 2), D. Field (3)

(1) Observatoire de Paris, Meudon, France

(2) Université de Cergy-Pontoise, Cergy Cedex

(3) Department of Physics and Astronomy, Århus University, Århus C, Denmark.

We present optical, near-infrared (NIR) images and spectra of the compact, dusty and complex HII region N88A in the SMC. The data were obtained mainly in the visible with the Boller & Chivens (B&C) spectrograph at the ESO 1.5m telescope and recently with the infrared spectrometer and array camera (ISAAC) at the VLT, around 2 μm. The NIR observations identified a region emitting in H_{2}, both in and around the two major components of N88. Using long-slit spectra obtained in the 3600-10100 Å wavelength range we have re-determined the key physical parameters of electron temperature and densities. The spectral type of the complex exciting source of N88A has been classified as an O6-O8 V, using He and H lines in the visible and NIR, with the HeI 1.7μm/2.11μm line ratio in the K band spectrum, suggesting that N88A is effectively ionized by a source of type < O7 V. However, from reddening correction and the temperature determined by photoionization mode l calculations, the ionizing source should also contain a very early type O star. A comparison with a recent photometric study shows that the V magnitude of the exciting source of N88A has previously been overestimated. The chemical abundances of He, O, N, Ne, S, and Ar, from emission-line intensities, are also derived. Abundances are consistent with the average abundances for SMC HII regions.

Comments: Accepted by: A&A e-mail: testor@mesiog.obspm.fr

Detection of Magnesium-Rich Ejecta in the Middle-Aged Supernova Remnant N49B

Sangwook Park (1), John P. Hughes (2), Patrick O. Slane (3), David N. Burrows (1), Jessica S. Warren (2), Gordon P. Garmire (1), and John A. Nousek (1)

(1) Pennsylvania State University, University Park, PA, USA

(2) Rutgers University, Piscataway, NJ, USA

(3) Harvard-Smithsonian Center for Astrophysics, Cambridge, MA, USA

The middle-aged supernova remnant (SNR) N49B in the Large Magellanic Cloud has been observed with the Chandra X-Ray Observatory. The superb angular resolution of Chandra resolves the complex structure of X-ray emitting filaments across the SNR. All observed features are soft (E ¡ 3 keV), and we find no evidence of either point-like or extended hard emission within the SNR. Spectral lines from O, Ne, Mg, Si, S, and Fe are present. Equivalent width images for the detected elemental species and spatially resolved spectral analysis reveal the presence of Mg-rich ejecta within the SNR. We find no such enrichment in O or Ne, which may reflect details of the nucleosynthesis process or the heating and cooling of the ejecta as it evolved. The bright circumferential filaments are emission from the shocked dense interstellar medium (ISM). We detect faint diffuse X-ray emission that extends beyond the X-ray bright filaments toward the west and southeast. These features appear to be the blast wave shock front expanding into lower density portions of the ISM seen in projection. We set an upper limit of 2 1033 ergs s-1 on the 0.55 keV band X-ray luminosity of any embedded compact object.

Comments: Published in ApJ Letters, July 30, 2003 (ApJ, 592, L41) WWW: http://www.journals.uchicago.edu/ApJ/journal/issues/ApJL/v592n1/17074/17074.html e-mail: park@astro.psu.edu

Massive Stars in the Local Group: Implications for Stellar Evolution and Star Formation

Philip Massey

Lowell Observatory, Flagstaff, AZ, USA

The galaxies of the Local Group serve as important laboratories for understanding the physics of massive stars. Here I discuss what is involved in identifying various kinds of massive stars in nearby galaxies: the hydrogen-burning O-type stars, and their evolved He-burning evolutionary descendants, the luminous blue variables, red supergiants, and Wolf-Rayet stars. Primarily I review what our knowledge of the massive star population in nearby galaxies has taught us about stellar evolution and star formation. I show that the current generation of stellar evolutionary models do well at matching some of the observed features, and provide a look at the sort of new observational data that will provide a benchmark against which new models can be evaluated.

Comments: To appear in Annual Reviews of Astronomy and Astrophysics, Vol 41, 15 WWW: ftp://ftp.lowell.edu/pub/massey/annualmassey.pdf e-mail: Phil.Massey@lowell.edu

The Optical Gravitational Lensing Experiment. Catalog of RR Lyrae Stars from the Small Magellanic Cloud

I. Soszynski (1,2), A. Udalski (1), M. Szymanski (1), M. Kubiak (1), G. Pietrzynski (1,3), P. Wozniak (4), K. Zebrun (1), O. Szewczyk (1), L. Wyrzykowski (1)

(1) Warsaw University Observatory, Warszawa, Poland

(2) Princeton University Observatory, Princeton, NJ, USA

(3) Universidad de Concepcion, Departamento de Fisica, Concepcion, Chile

(4) Los Alamos National Laboratory, Los Alamos, NM, USA

We present the catalog of RR Lyrae stars from 2.4 square degrees of central parts of the Small Magellanic Cloud (SMC). The photometric data were collected during four years of the OGLE-II microlensing survey. Photometry of each star was obtained using the Difference Image Analysis (DIA) method. The catalog contains 571 objects, including 458 RRab, 56 RRc variables, and 57 double mode RR Lyr stars (RRd). Additionally we attach a list of variables with periods between 0.18-0.26 days, which are probably delta Sct stars. Period, BVI photometry, astrometry, amplitude, and parameters of the Fourier decomposition of the I-band light curve are provided for each object. We also present the Petersen diagram for double mode pulsators.

We found that the SMC RR Lyr stars are fairly uniformly distributed over the studied area, with no clear correlation to any location. The most preferred periods for RRab and RRc stars are 0.589 and 0.357 days, respectively. We noticed significant excess of stars with periods of about 0.29 days, which might be second-overtone RR Lyr stars (RRe). The mean extinction free magnitudes derived for RRab stars are 18.97, 19.45 and 19.73 mag for the I, V and B-band, respectively.

All presented data, including individual BVI observations, are available from the OGLE Internet archive.

Comments: Acta Astronomica (2002) 52, 369 WWW: http://arXiv.org/abs/astro-ph/0212410 e-mail: soszynsk@astrouw.edu.pl

The Optical Gravitational Lensing Experiment. Catalog of RR Lyrae Stars in the Large Magellanic Cloud

I. Soszynski (1), A. Udalski (1), M. Szymanski (1), M. Kubiak (1), G. Pietrzynski (1,2), P. Wozniak (3), K. Zebrun (1), O. Szewczyk (1), L. Wyrzykowski (1)

(1) Warsaw University Observatory, Warszawa, Poland

(2) Princeton University Observatory, Princeton, NJ, USA

(3) Universidad de Concepcion, Departamento de Fisica, Concepcion, Chile

(4) Los Alamos National Laboratory, Los Alamos, NM, USA

We present the catalog of RR Lyr stars discovered in a 4.5 square degrees area in the central parts of the Large Magellanic Cloud (LMC). Presented sample contains 7612 objects, including 5455 fundamental mode pulsators (RRab), 1655 first-overtone (RRc), 272 second-overtone (RRe) and 230 double-mode RR Lyr stars (RRd). Additionally we attach a list of several dozen other short-period pulsating variables. The catalog data include astrometry, periods, BVI photometry, amplitudes, and parameters of the Fourier decomposition of the I-band light curve of each object.

We present density map of RR Lyr stars in the observed fields which shows that the variables are strongly concentrated toward the LMC center. The modal values of the period distribution for RRab, RRc and RRe stars are 0.573, 0.339 and 0.276 days, respectively. The period-luminosity diagrams for BVI magnitudes and for extinction insensitive index W_{I} are constructed. We provide the logP-I, logP-V and logP-W_{I} relations for RRab, RRc and RRe stars. The mean observed V-band magnitudes of RR Lyr stars in the LMC are 19.36 mag and 19.31 mag for ab and c types, respectively, while the extinction free values are 18.91 mag and 18.89 mag.

We found a large number of RR Lyr stars pulsating in two modes closely spaced in the power spectrum. These stars are believed to exhibit non-radial pulsating modes. We discovered three stars which simultaneously reveal RR Lyr-type and eclipsing-type variability. If any of these objects were an eclipsing binary system containing RR Lyr star, then for the first time the direct determination of the mass of RR Lyr variable would be possible.

We provide a list of six LMC star clusters which contain RR Lyr stars. The richest cluster, NGC 1835, hosts 84 RR Lyr variables. The period distribution of these stars suggests that NGC 1835 shares features of Oosterhoff type I and type II groups.

All presented data, including individual BVI observations and finding charts are available from the OGLE Internet archive. Comments: Acta Astronomica 53 (2003) 93 WWW: http://arXiv.org/abs/astro-ph/0306041 e-mail: soszynsk@astrouw.edu.pl

Cluster Mass Functions in the Large and Small Magellanic Clouds: Fading and Size of Sample Effects

D. A. Hunter (1), B. G. Elmegreen (2), T. J. Dupuy (1), and M. Mortonson (1)

(1) Lowell Observatory, Flagstaff, AZ, USA

(2) IBM T.J. Watson Research Center, Yorktown Heights, NY, USA

The properties of 939 star clusters in the Large and Small Magellanic Clouds were determined from ground-based CCD images in UBVR passbands. The area coverage was extensive, corresponding to 11.0 sq-kpc in the LMC and 8.3 sq-kpc in the SMC. After corrections for reddening, the colors and magnitudes of the clusters were converted to ages and masses, and the resulting mass distributions were searched for the effects of fading, evaporation, and size-of-sample bias. The data show a clear signature of cluster fading below the detection threshold. The initial cluster mass function (ICMF) was determined by fitting the mass and age distributions with cluster population models. These models suggest a new method to determine the ICMF that is nearly independent of fading or disruption and is based on the slope of a correlation between age and the maximum cluster mass in equally spaced intervals of log-age. For a nearly uniform star formation rate, this correlation has a slope equal to 1/( α-1) for an ICMF of dn(M)/dM proportional to M^{−α}. We determine that alpha is between 2 and 2.4 for the LMC and SMC using this method plus another method in which models are fit to the mass distribution integrated over age and to the age distribution integrated over mass. The maximum mass method also suggests that the cluster formation rate in the LMC age gap between 3 and 13 Gy is about a factor of ten below that in the period from 0.1 Gy to 1 Gy. The oldest clusters correspond in age and mass to halo globular clusters in the Milky Way. They do not fit the trends for lower-mass clusters but appear to be a separate population that either had a very high star formation rate and became depleted by evaporation or formed with only high masses.

Comments: To appear in AJ, October 2003. WWW: ftp://ftp.lowell.edu/pub/dah/papers/mccl e-mail: dah@lowell.edu

A theoretical analysis of the systematic errors in the Red Clump distance to the LMC

Maurizio Salaris (1), Susan Percival (1), Leo Girardi (2)

(1) Astrophysics Research Institute, Liverpool John Moores University, Birkenhead, UK

(2) INAF – Osservatorio Astronomico di Trieste, Trieste, Italy

We present a detailed analysis of the uncertainty on the theoretical population corrections to the LMC Red Clump (RC) absolute magnitude, by employing a population synthesis algorithm to simulate theoretically the photometric and spectroscopic properties of RC stars, under various assumptions about the LMC Star Formation Rate (SFR) and Age Metallicity Relationship (AMR). A comparison of the outcome of our simulations with observations of evolved low-intermediate mass stars in the LMC allows one to select the combinations of SFR and AMR that bracket the real LMC star formation history, and to estimate the systematic error on the associated RC population corrections. The most accurate estimate of the LMC distance modulus from the RC method (adopting the OGLE-II reddening maps for the LMC) is obtained from the K-band magnitude, and provides (m-M) _{(0,LMC)}=18.47+-0.01(random) _{(−0.06)}^{(+0.05)} (systematic). Distances obtained from the I-band, or from the multicolour RC technique which determines at the same time reddening and distance, both agree (albeit with a slightly larger error bar) with this value.

Comments: MNRAS, in press WWW: http://xxx.soton.ac.uk/abs/astro-ph/0307329 e-mail: ms@astro.livjm.ac.uk

Abstracts of Non-Refereed Papers

An Application of Bayesian Statistics to the Extragalactic Cepheid Distance Scale

Thomas G. Barnes III (1), W. H. Jefferys (1), Thomas J. Moffett (2), and Amy D. Forestell (1)

(1) University of Texas, Austin, TX, USA

(2) Purdue University, West Lafayette, IN, USA

We have determined quasi-geometric distances to the Magellanic Clouds. Our analysis uses the Bayesian statistical method of Barnes et al. (2003, ApJ, 592, July 20; astro-ph/0303656) to provide mathematically rigorous and objective solutions to individual Cepheids. We combine the individual Cepheid distances with a hierarchical Bayesian model to determine the galactic distances. We obtain distance moduli 18.87 0.07 mag (LMC, 12 stars) and 19.14 0.10 (SMC, 8 stars) - uncorrected for metallicity. If the PL relations of the LMC, SMC and Galaxy are identical, our results exclude the metallicity effect predicted by Hindsley & Bell (1989, ApJ, 341, 1004) in the V, (V-R) surface brightness method at the 5 sigma level. Alternately, if Hindsley & Bell’s prediction is adopted as true, we find a metallicity effect intrinsic to the Cepheid PL relation requiring a correction delta(V-Mv) = (0.36 0.07) delta[A/H] mag. The latter has the opposite sign to other observational estimates of the Cepheid metallicity effect. The radii we determine are larger by 0.05 in the log than a Galactic period-radius relation determined by classical methods. Resolution of this offset will be addressed in a later paper.

Comments: Presented at IAU Colloquium 193 ”Variable Stars in the Local Group,” Christchurch, New Zealand, July 7-11, 2003. WWW: not yet available e-mail: tgb@astro.as.utexas.edu

Recent astro-ph Listings

astro-ph/0307509:
      Title: H_alpha Emission from High-Velocity Clouds and their Distances
      Authors: M.E. Putman (U. Colorado), J. Bland-Hawthorn (AAO), S. Veilleux (U.
      Maryland), B.K. Gibson (Swinburne), K.C. Freeman (RSAA), P.R. Maloney (U. Colorado)
      Comments: ApJ accepted, see this http URL for the full resolution version
astro-ph/0307431:
      Title: Using Optical and Near-IR Photometry to Test MACHO Lens Candidates
      Authors: Ted von Hippel, Ata Sarajedini, Maria Teresa Ruiz
      Comments: 28 pages, 8 figures, ApJ, accepted for Oct 1 issue

astro-ph/0307419:
Title: XMM-Newton RGS spectroscopy of LMC X-3
Authors: M.J. Page (MSSL), R. Soria (MSSL), K. Wu (MSSL), K.O. Mason (MSSL), F.A.
Cordova (UCR), W.C. Priedhorsky (LANL)
Comments: Accepted for publication in MNRAS
astro-ph/0307375:
Title: Gamma-ray and synchrotron emission from neutralino annihilation in the Large
Magellanic Cloud
Authors: Argyro Tasitsiomi, Jennifer Gaskins, Angela V. Olinto
Comments: 10 (twocolumn) pages, 3 figures
astro-ph/0307373:
Title: Correlated Optical and X-ray Variability in LMC X-2
Authors: Katherine E. McGowan (1), Phil A. Charles (2), Darragh O'Donoghue (3), Alan
P. Smale (4) ((1) LANL, (2) University of Southampton, (3) SAAO, (4) NASA/GSFC)
Comments: 11 pages, 9 figures, accepted for publication in MNRAS
astro-ph/0307351:
Title: Space Telescope Imaging Spectrograph slitless observations of Small Magellanic
Cloud Planetary Nebulae: a study on morphology, emission line intensity, and evolution
Authors: Letizia Stanghellini (STScI), Richard A. Shaw (NOAO), Bruce Balick (U.
Washington), Max Mutchler (STScI), J. Chris Blades (STScI), Eva Villaver (STScI)
Comments: Accepted for publication in the Astrophysical Journal, 30 pages, 13 figures,
6 tables. For high resolution version of Figs 1 to 6, see this http URL
astro-ph/0307329:
Title: A theoretical analysis of the systematic errors in the Red Clump distance to the
LMC
Authors: Maurizio Salaris (1), Susan Percival (1), Leo Girardi (2) ((1) ARI, Liverpool John
Moores University, (2) INAF - Osservatorio Astronomico di Trieste)
Comments: Accepted for publication in MNRAS
astro-ph/0307250:
Title: Post-AGB Evolution in the Large Magellanic Cloud. A Study of the Central Stars of
Planetary Nebulae
Authors: Eva Villaver (STScI), Letizia Stanghellini (STScI), Richard A. Shaw (NOAO)
Comments: Accepted for publication in Astrophysical Journal. 23 pages, 8 figures and 4
tables
astro-ph/0307145:
Title: A Quest for PMS candidate stars at low metallicity: Variable HAe/Be and Be stars
in the Small Magellanic Cloud
Authors: W.J. de Wit (1), J-P. Beaulieu (2), H.J.G.L.M. Lamers (3,4), E. Lesquoy (2), J-B.
Marquette (2); ((1) Arcetri (Florence), (2) IAP (Paris), (3) University of Utrecht (Utrecht),
(4) SRON (Utrecht))
Comments: 18pages, 22 figures, accepted for publication in A&A
astro-ph/0307016:
Title: Distances and ages of NGC 6397, NGC 6752 and 47 Tuc
Authors: Raffaele Gratton (1), Angela Bragaglia (2), Eugenio Carretta (1), Gisella
Clementini (2), Silvano Desidera (1), Frank Grundahl (3), Sara Lucatello (1,4) ((1)
INAF-Osservatorio Astronomico di Padova, ITALY, (2) INAF- Osservatorio Astronomico
di Bologna, ITALY, (3) Aarhus Univ., DENMARK, (4) Dip. di Astronomia, Univ. Padova,
ITALY)
Comments: 16 pages, 5 encapsulated figures, accepted for publication in A&A

astro-ph/0306561:
Title: Neutralino annihilation gamma-rays from clumps and the LMC
Authors: Argyro Tasitsiomi, Jennifer Gaskins, Angela V. Olinto
Comments: 4 pages, 4 figures; to appear in New Astronomy Reviews, Proceedings of
2nd VERITAS symposium
astro-ph/0306528:
Title: Cluster Mass Functions in the Large and Small Magellanic Clouds: Fading and Size
of Sample Effects
Authors: D. A. Hunter (1), B. G. Elmegreen (2), T. J. Dupuy (1), M. Mortonson (1) ((1)
Lowell Observatory, (2) IBM T. J. Watson Research Center)
Comments: To be published in AJ, October 2003. Available from ftp.lowell.edu, cd
pub/dah/papers/mccl
astro-ph/0306501:
Title: On the Structure and Kinematics of Nebulae around LBVs and LBV Candidates in
the LMC
Authors: Kerstin Weis (ITA, Heidelberg; UMN, Minnesota; MPIfR Bonn)
Comments: paper accepted by A&A, 25 pages, 24 figures; paper with images in full
resolution available at this http URL
astro-ph/0306426:
Title: Red variables in the OGLE-II database. I. Pulsations and period-luminosity
relations below the tip of the Red Giant Branch of the LMC
Authors: L.L. Kiss, T. Bedding
Comments: 5 pages, 4 figures; accepted for publication in MNRAS (Pink Pages); proof
corrections added
astro-ph/0306425:
Title: The Spectrum of Integrated Millimeter Flux of the Magellanic Clouds and
30-Doradus from TopHat and DIRBE Data
Authors: J. E. Aguirre, J. J. Bezaire, E. S. Cheng, D. A. Cottingham, S. S. Cordone, T. M.
Crawford, D. J. Fixsen, L. Knox, S. S. Meyer, H. U. Norgaard-Nielsen, R. F. Silverberg, P.
Timbie, G. W. Wilson
Comments: 41 pages, 4 figures. Accepted for publication in Astrophysical Journal
astro-ph/0306358:
Title: Magellanic Cloud Star Clusters as Simple Stellar Populations
Authors: Andrew. J. Leonardi, James. A. Rose
Comments: 29 pages, 25 figures, 7 tables, accepted for publication in the Astronomical
Journal
astro-ph/0306318:
Title: Detection of Magnesium-Rich Ejecta in the Middle-Aged Supernova Remnant
N49B
Authors: Sangwook Park, John P. Hughes, Patrick O. Slane, David N. Burrows, Jess ica
S. Warren, Gordon P. Garmire, John A. Nousek
Comments: 3 text pages (ApJ emulator style), 3 figures, 1 table, Accepted for the
publication in Ap J Letters
Journal-ref: Astrophys.J. 592 (2003) L41-L44
astro-ph/0306144:
Title: Unusual CO Line Ratios and kinematics in the N83/N84 Region of the Small
Magellanic Cloud
Authors: Alberto D. Bolatto, Adam Leroy (UC Berkeley), James M. Jackson (Boston U.),
Frank P. Israel (Leiden Obs.)
Comments: 29 pages, aastex preprint style, 7 B&W figures (some degraded to fit the file
size constraints), accepted for publication in ApJ, retrieve paper with full resolution
figures from this http URL

astro-ph/0306116:
Title: XTE J0111.2-7317 : a nebula-embedded X-ray binary in the SMC
Authors: M.J. Coe (Southampton University), N.J. Haigh (Southampton University), C.A.
Wilson (NASA/MSFC), I. Negueruela (Alicante University)
Comments: Accepted for publication by MNRAS
astro-ph/0306084:
Title: WR bubbles and HeII emission
Authors: Y. Naze (1), G. Rauw (1), J. Manfroid (1), Y.-H. Chu (2), J.-M. Vreux (1) ((1)
Institut d'Astrophysique et de Geophysique, Liege, Belgium; (2) University of Illinois at
Urbana-Champaign, USA)
Comments: 13 pages, 8 figures (7 in jpg), accepted by A&A, also available from this http
URL
astro-ph/0306041:
Title: The Optical Gravitational Lensing Experiment. Catalog of RR Lyrae Stars in the
Large Magellanic Cloud
Authors: I. Soszynski, A. Udalski, M. Szymanski, M. Kubiak, G. Pietrzynski, P. Wozniak, K.
Zebrun, O. Szewczyk, L. Wyrzykowski
Comments: 25 pages, 10 figures, LaTeX. Figures 1, 3, 6, 7, 8, 9 included in GIF format.
Photometric data presented in the paper are available from the OGLE Internet archive:
this http URL or its US mirror this http URL
Journal-ref: Acta Astron. 53 (2003) 93
astro-ph/0306038:
Title: Results of the ESO-SEST Key Programme on CO in the Magellanic Clouds X. CO
emission from star formation regions in LMC and SMC
Authors: F.P. Israel (Sterrewacht Leiden, NL), L.E.B. Johansson (Onsala, Sweden), M.
Rubio, G. Garay (Dpto de Astronomia, U. de Chile), Th. de Graauw (SRON Groningen,
NL), R.S. Booth (Onsala, Sweden), F. Boulanger (ENS and IAS, Paris, Fr), M.L. Kutner
(Astron. Dpt, UoTexas, USA), J. Lequeux (Obs. de Paris, Fr.), L.-A. Nyman (Onsala,
Sweden; ESO, Chile)
Comments: 13 pages, accepted by Astronomy and Astrophysics
astro-ph/0305574:
Title: Discovery of a New Pulsar Wind Nebula in the Large Magellanic Cloud
Authors: B. M. Gaensler (1), S. P. Hendrick (2), S. P. Reynolds (1, 2), K. J. Borkowski (2)
((1) Harvard-Smithsonian CfA, (2) NCSU)
Comments: 4 pages of text plus 2 embedded EPS figures. Minor changes following
referee's report. ApJ Letters, in press